Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

cards

Expert replies
by orel » Wed Mar 25, 2009 6:35 pm
...
You do not have the required permissions to view the files attached to this post.
Join the discussion
Source: — Problem Solving |

by krisraam » Wed Mar 25, 2009 7:21 pm
Finding the probability of selecting cards that are not pairs.

Probability of selecting first card = 12/12 = 1 ( It can be any card).

Probability of selecting second card = 10/11 ( 11 cards are left for selection. We shouldn't select the matching card. )

Probability of selecting the third card = 8/10 ( 10 cards left we shouldn't select the two cards that match the already selected 2 cards)

Probability of selecting the 4th card = 6/9

Probability of selecting the cards with no pairs = 1* 10/11*8/10*6/9 = 16/33

Probability of selecting at least one pair = 1-16/33 = 17/33.

Thanks
Raama
Join the discussion

by vittalgmat » Thu Mar 26, 2009 12:16 am
On an slightly related topic. (sorry to post here. I will mail eric and mozilla guys about this issue)
Does everyone see jpg attachments properly in firefox for windows. The community sponsors banner drops down over the picture and I cant read completely.

thanks
-V
Join the discussion

by orel » Thu Mar 26, 2009 12:52 am
Vittal,

the same happens to me. I cannot properly see the attachments on posts. :(
Join the discussion

by avenus » Thu Mar 26, 2009 1:47 am
On an slightly related topic. Does everyone see jpg attachments properly in firefox for windows. The community sponsors banner drops down over the picture and I cant read completely. thanks -V
A quick solution: place the mouse pointer over the image, click on the right button and select show image or the equivalent in the language/browser you're using. The image on its own will be shown in a separate tab...
Join the discussion

by kanha81 » Thu Mar 26, 2009 9:15 am
krisraam wrote:Finding the probability of selecting cards that are not pairs.

Probability of selecting first card = 12/12 = 1 ( It can be any card).

Probability of selecting second card = 10/11 ( 11 cards are left for selection. We shouldn't select the matching card. )

Probability of selecting the third card = 8/10 ( 10 cards left we shouldn't select the two cards that match the already selected 2 cards)

Probability of selecting the 4th card = 6/9

Probability of selecting the cards with no pairs = 1* 10/11*8/10*6/9 = 16/33

Probability of selecting at least one pair = 1-16/33 = 17/33.

Thanks
Raama
I really like the way you solve the problems, makes it look simpler and understandable. I always fight between 2 methods. This is the method that I chose; however, it took more than 2 mins, but less than 4 mins.

P(E): Finding at least 1 pair that have same value.
4 cards chosen from 12: 12C4 ways

We find P(E`): Finding none of the 4 cards chosen have same value
=6C4 * 4C1 * 4 (ie choose 4 cards from a pair of 6 cards, such that, you choose 1 card from each pair and this can be done in 4 ways) / (12C4)

= 16/33

p(E) = 1-P(E`) = 1 - 16/33 = 17/33
Want to Beat GMAT.
Always do what you're afraid to do. Whoooop GMAT
Join the discussion

by vittalgmat » Thu Mar 26, 2009 2:42 pm
avenus wrote:
On an slightly related topic. Does everyone see jpg attachments properly in firefox for windows. The community sponsors banner drops down over the picture and I cant read completely. thanks -V
A quick solution: place the mouse pointer over the image, click on the right button and select show image or the equivalent in the language/browser you're using. The image on its own will be shown in a separate tab...
Thanks avenus. As u mentioned, I right clicked on the Ad and chose "block images". This blocked images from page2ad.google.com.
Join the discussion

by RBurroughs » Thu Mar 26, 2009 8:05 pm
vittalgmat wrote:On an slightly related topic. (sorry to post here. I will mail eric and mozilla guys about this issue)
Does everyone see jpg attachments properly in firefox for windows. The community sponsors banner drops down over the picture and I cant read completely.

thanks
-V
you can also right click and select "do not show images from this site" to block advertisements
Join the discussion

by Vemuri » Fri Mar 27, 2009 5:48 am
krisraam wrote:Finding the probability of selecting cards that are not pairs.

Probability of selecting first card = 12/12 = 1 ( It can be any card).

Probability of selecting second card = 10/11 ( 11 cards are left for selection. We shouldn't select the matching card. )

Probability of selecting the third card = 8/10 ( 10 cards left we shouldn't select the two cards that match the already selected 2 cards)

Probability of selecting the 4th card = 6/9

Probability of selecting the cards with no pairs = 1* 10/11*8/10*6/9 = 16/33

Probability of selecting at least one pair = 1-16/33 = 17/33.

Thanks
Raama
I am confused here....appreciate some help clearing out the cloud for me. Thanks in advance.

When I look at the way the question is framed, I am considering that 4 cards have been laid out & we have 8 cards remaining to find the probability of not finding a match. Can someone please help with explaining the approach.
-------------------------------------------------
Earth Hour 2009
Vote Earth! Your light switch is your vote.
https://www.earthhour.org/home/
Join the discussion

by rs2010 » Fri Mar 27, 2009 9:08 am
vittalgmat wrote:On an slightly related topic. (sorry to post here. I will mail eric and mozilla guys about this issue)
Does everyone see jpg attachments properly in firefox for windows. The community sponsors banner drops down over the picture and I cant read completely.

thanks
-V
Safari, Chrome and Firefox also. Think about Mac OS, which does not have IE.

Work around: to view properly right click on image and view it in another tab.

I feel its more with formatting than with browsers.
Join the discussion